OpenStudy (anonymous):

Find equations of both the tangent lines to the ellipse x2 + 4y2 = 36 that pass through the point (12, 3).

OpenStudy (amistre64):

might want to develop the derivative of the ellipse for a any general point

OpenStudy (amistre64):

can you derive the ellipse using implicit differentation?

OpenStudy (anonymous):

wouldn't it just be 2x+8y=0?

OpenStudy (amistre64):

yes, but with a y' added in for good measure

OpenStudy (amistre64):

2x+8yy'=0 y'=-x/y for any (x,y) on the ellipse

OpenStudy (amistre64):

deleted me 4

OpenStudy (amistre64):

y'=-x/4y

OpenStudy (anonymous):

so now I just plug in the point(12,3) to get the slop for the tangent equation?

OpenStudy (anonymous):

slope**

OpenStudy (amistre64):

so lets take a general point (a,b) along the ellipse to define its slope y' = -a/4b also has to be the slope of the line that goes thru (12,3) \[y=-\frac{a}{4b}(x-12)+3\]but (a,b) also on that line \[b=-\frac{a}{4b}(a-12)+3\] \[b=-\frac{a^2}{4b}-\frac{12}{4b}+3\] \[4b^2=-a^2+12+12b\] \[a^2+4b^2-12b=12\] \[a^2+4(b^2-3b)=12\] \[a^2+4(b^2-3b+\frac94-\frac94)=12\] \[a^2+4(b-3)^2-9=12\] \[a^2+4(b-3)^2=21\] \[\frac{a^2}{21}+\frac{4(b-3)^2}{21}=1\] make sure i did those right

OpenStudy (amistre64):

essentially, we create another ellipse; where this ellipse intersects the original ellipse will define the 2 points that are needed

OpenStudy (amistre64):

i see one error, that doesnt really change things to much\[(b-\frac32)^2\]

OpenStudy (amistre64):



OpenStudy (anonymous):

wouldn't it be b=−a2/4b+12a/4b+3?

OpenStudy (anonymous):

and then we solve from there? (this is near the very top)

OpenStudy (amistre64):

yes, then the last term becomes 3*4b/4b = 12b/4b and we can collect all the numberators into one place; cross multiply the 4b over; and simplify

OpenStudy (amistre64):

i fixed the typo in the next line :)

OpenStudy (anonymous):

oh, that makes sense. Okay, I think I understand it for the most part.

OpenStudy (amistre64):

its fairly simple, it just making sure you dont trip over te algebra along the way :)

OpenStudy (anonymous):

so I understand all the algebra but I am unsure on how to find the exact equations of the tangent lines... can you help me on that?

OpenStudy (amistre64):

well, to see what it is graphically http://www.wolframalpha.com/input/?i=y%3D-x%2F%284y%29%28x-12%29%2B3%2C+x%5E2+%2B+4y%5E2+-+36+%3D+0

OpenStudy (amistre64):

set the equations of the ellipses equal to each other; im not sure that i typed out the process correctly tho ...

OpenStudy (amistre64):

\[b=-\frac{a}{4b}(a-12)+3\] \[4b^2=-a(a-12)+12b\] \[4b^2-12b=-(a^2-12a)\] \[4(b^2-3b+\frac94)-9=-(a^2-12a+36)+36\] \[4(b-\frac32)^2-9=-(a-6)^2+36\] \[(a-6)^2+4(b-\frac32)^2=45\] thats better

OpenStudy (anonymous):

that makes sense :) so how do I get from there to the equations? I am lost :/

OpenStudy (anonymous):

I am confused on how to find the equations of the tangent lines